# The role of antisymmetric orbitals and electron-electron interactions on the two-particle spin and valley blockade in graphene double quantum dots

### v1

@@ Abstract
To electrostatically define a DQD, we follow previous works[22, 23, 24, 25, 50, 51, 26] and fabricate heterostructures consisting of a BLG crystal encapsulated within two layers of hexagonal boron nitride (with a thickness of $\approx 30$ nm each) [52, 53]. The heterostructure is placed on a graphite flake which acts as a back gate (BG) [54]. Split gates (SGs) consisting of $\approx 25$ nm Cr/Au are deposited on top, forming a $\approx 100$ nm wide channel. An additional $\approx 75$ nm thick layer of metallic Cr/Au finger gates (FGs) is then fabricated across the channel. The top gate layers are separated by a 15 nm thick layer of atomic layer deposited Al2O3. A schematic top view of a finished sample is shown in Fig. 1(a). All measurements are perfomed in a dilution refrigerator with a base temperature below 30 mK. Voltages of opposite sign applied to the SGs and BG create an electric displacement field perpendicular to the plane of the BLG, opening a band gap and allowing the Fermi energy to be tuned into the gap [54, 35]. This leaves a narrow conductive channel, defined by the SGs, connecting the source and drain reservoirs. Voltages applied to two adjacent FGs ($V_{\mathrm{L}},V_{\mathrm{R}}$) are used to locally invert the polarity of the channel, creating a DQD and giving rise to tunnel barriers where the Fermi energy lies within the band gap, as illustrated in Fig. 1(b). At a finite bias voltage ($V_{\mathrm{SD}}$), we measure the conduction through the channel as a function of $V_{\mathrm{L}}$ and $V_{\mathrm{R}}$, as shown in Fig. 1(c). For $V_{\mathrm{L}}>4.2$ V and $V_{\mathrm{R}}>4.05$ V, we observe the formation of an electron-electron DQD  [^1] For lower voltages, a hole QD is confined in between the two FGs, which can form an ambipolar electron-hole DQD, see also Refs. [73, 25, 28], which is incrementally filled for increasingly positive FG voltages. White numbers $(N_{\mathrm{L}},N_{\mathrm{R}})$, indicate the charge occupation in the left and right QD, respectively.
@@
To understand the observed resonant features and the blockade effects, we discuss possible transitions between single- and two-particle states in a DQD when including both symmetric and antisymmetric orbital states. For the investigated energy range we focus on 16 relevant two-particle states, that group into 6 orbitally symmetric and 10 orbitally antisymmetric states, separated by the orbital splitting $\Delta_{\mathrm{Orb}}$ [57, 58]. The orbitally symmetric states $\phi_{s}$ are further split by short-range electron-electron interactions into a valley singlet – spin triplet ground state (GS) and valley triplet – spin singlet excited states (ESs) with a separation of $\delta_{2}$ [57, 59, 58, 60, 61], as shown in Fig. 3(a). The next higher orbital state is antisymmetric $\phi_{a}$, featuring spin triplet - valley triplet states and a spin singlet - valley singlet state, which are separated by the Kane-Mele spin-orbit coupling $\Delta_{\mathrm{SO}}\approx 65~{}\mu$eV [62, 26, 25]. For the $(1,1)$ states, we assume a negligible mixing between the two QDs due to a small interdot tunnel coupling and treat them as independent single particle states in each QD.
First, we focus at the $(1,1)\rightarrow(2,0)$ charge transition at $V_{\mathrm{SD}}<0$ V. For simplicity, we initially only consider transitions from the $(1,1)$ GS to the 16 different two-particle states. Fig. 3(a) displays possible transitions from the $(1,1)$ GS, which is the |$\ket{K^{-}\uparrow,\,K^{-}\uparrow}$ state for $B_{\perp}\geq 0$ [62, 26, 25], into spin triplet - valley singlet (green), spin singlet - valley triplet (blue) and spin triplet - valley triplet states (red) of the symmetric and antisymmetric two-particle states in the $(2,0)$ configuration at $B_{\perp}=0$. Increasing the perpendicular magnetic field shifts all states according to their spin and valley magnetic moments [57], which is illustrated in Fig. 3(b). The condition for a charge transition to resonantly occur is that the detuning energy of the DQD needs to compensate the energy difference between the electrochemical potentials of the $(1,1)$ and $(2,0)$ states, given by
@@
with the Bohr magneton, $\mu_{B}$, and the valley g-factor of the single particle orbital, $g^{(1)}_{v}$. For $B>B_{\mathrm{TS}}$, where the two-particle GS changes to |$\ket{S^{s}T^{v}_{-}}$, the GS to GS transition is |$\ket{K^{-}\uparrow,\,K^{-}\uparrow}\rightarrow\ket{S^{s}T^{v}_{-}}$, which shifts as
@@
Importantly, Fig. 3(c) illustrates that we expect to observe a magnetic field tunable valley and spin blockade in the $(1,1)\rightarrow(2,0)$ bias triangle, as indicated by the green and blue colored regions, respectively. For $B_{\perp}<B_{\mathrm{TS}}$, the $(1,1)$ GS to $(2,0)$ GS transition |$\ket{K^{-}\uparrow,\,K^{-}\uparrow}\rightarrow\ket{T^{s}_{-}S^{v}}$ (green region) requires a valley flip. Valley blockade (VB) is therefore only lifted if the antisymmetric orbital states are accessible, limiting the detuning energy range of the valley blocked region to
@@
The position of the zero detuning energy in the gate space changes slightly during the measurement, firstly due to slow electrostatic noise in the sample, and secondly due to the GS to GS transition shifting with the magnetic field. Therefore, using magnetic field dependent measurements of the bias triangles (see Fig. 2), the detuning traces are shifted such that $\varepsilon=0$ is always at the same position, as indicated by the white dashed line in Fig. 4(a). Our simulation is able to reproduce many features in the data using $\Delta_{\mathrm{Orb}}=0.7$ meV, $\delta_{2}=0.34$ meV, $g^{(1)}_{v}=15$, $g^{s}_{v}=18$, $g^{a}_{v}=19$, which is in agreement with other measurements on QDs of similar dimension [37, 25, 26]. Both simulation and measurements show the transition from valley blockade (marked by the star) to spin blockade (marked by the circle), as can be seen by comparing Figs. 4(a) and 4(b). We identify prominent resonances labeled a - d in the data  [^2] The software of the simulation allows to interactively inspect any position in gate space, where it illustrates the alignment of the chemical potentials in the left and right QD., with resonance a corresponding to a transition from the $(1,1)$ GS to the valley unpolarized triplet state in the antisymmetric orbital, |$\ket{K^{-}\uparrow,\,K^{-}\uparrow}\rightarrow\ket{T^{s}_{-}T^{v}_{0}}$, and requiring a valley flip. Resonance b corresponds to the transition that first lifts the valley and spin blockade, |$\ket{K^{-}\uparrow,\,K^{-}\uparrow}\rightarrow\ket{T^{s}_{-}T^{v}_{-}}$, and is strongly pronounced in both data and simulation. Resonance c involves a $(1,1)$ ES and requires a valley flip, |$\ket{K^{-}\uparrow,\,K^{-}\downarrow}\rightarrow\ket{S^{s}T^{v}_{0}}$. Finally, different valley blocked transitions contribute to resonance d, |$\ket{K^{-}\uparrow,\,K^{+}\downarrow}\rightarrow\ket{T^{s}_{0}T^{v}_{-}}$, |$\ket{K^{-}\downarrow,\,K^{+}\downarrow}\rightarrow\ket{T^{s}_{+}T^{v}_{-}}$, |$\ket{K^{-}\downarrow,\,K^{+}\uparrow}\rightarrow\ket{T^{s}_{0}T^{v}_{-}}$, which is therefore rather broadened. Some of the resonant features do not shift linearly with $B_{\perp}$, which has been observed before in similar measurements and can be caused by the magnetic field slightly changing the effective confinement potential and thereby changing the wavefunction dependent valley g-factors. Additionally, we observe a dip in conductance at $B_{\perp}\approx\pm 0.05$ and 0.10.10.1~{}0.1T. Similar effects can be observed in Fig. 10 and in equivalent measurements of BLG DQDs in the $(1,2)\leftrightarrow(0,3)$ charge configuration reported in Ref. [34], but have not yet been understood comprehensively.
@@
Following the same procedure as presented in the context of Fig. 3(c), we are able to plot the $B_{\perp}$ dependency of the expected resonant features as a function of detuning and perpendicular magnetic field, shown in Fig. 5(a). Here, the change in the two particle GS does not affect the detuning dependency of the transitions, because it also affects the onset of the bias triangle (defining $\varepsilon=0$), see also equation (The role of antisymmetric orbitals and electron-electron interactions on the two-particle spin and valley blockade in graphene double quantum dots). Therefore, Fig. 5(a) basically shows the magnetic field dispersion of the $(1,1)$ states with the energetically lowest state set to zero. For $B_{\mathrm{TS}}>B_{\perp}>0$, we expect an increasingly large valley blocked region (green region in Fig. 5(a)) since the two particle GS is a valley singlet, |$\ket{T^{s}_{-}S^{v}}$, while the $(1,1)$ GS is valley polarized. The blockade is lifted as soon as the valley unpolarized $(1,1)$ states are available, which occurs at
@@
For $B_{\perp}>B_{\mathrm{TS}}$, the two-particle GS becomes a spin singlet - valley triplet, |$\ket{S^{s}T^{v}_{-}}$, immediately lifting the valley blockade. Instead, a small spin blockade (blue region in Fig. 5(a)) appears close to zero detuning since the $(1,1)$ GS is spin polarized. Black arrows in Fig. 5(a) indicate the spin orientation of the three lowest $(1,1)$ states. The spin blockade is lifted as soon as the first spin unpolarized state is available
@@
We experimentally probe the $(2,0)\rightarrow(1,1)$ charge transition by measuring magnetic field dependent detuning cuts through the bias triangle, similar to Fig. 4(a), but for positive bias. These are presented in Fig. 5(b). Using the same parameters as for negative bias direction (see Fig. 4(b)), our simulation is able to reproduce the main features of the experimental data. Interestingly, both data and simulation show that the valley blocked region (star) extends beyond $B_{\mathrm{TS}}\approx 0.4$ T, which is in contrast to Fig. 5(a). Furthermore, the simulation predicts that the spin blocked region (circle) appears along the entire magnetic field range, not only for $B>B_{\mathrm{TS}}$. The spin blocked region predicted by the simulation is barely visible in the data, it can only be surmised close to zero detuning for $B_{\perp}\approx 0.05-0.15$ T. However, this can be caused by difficulties in estimating the onset of the bias triangle due to significant co-tunneling, see white dashed lines in Fig. 2(b).
@@ I.2 Rate equation based transport simulation
The software further allows to interactively inspect the charge stability diagrams, as illustrated by Fig. 11, where the DQD parameters in the simulation were set to $\Delta_{\mathrm{Orb}}=0.575$ meV, $\delta_{2}=0.2$ meV, $\delta_{1}=0$, $g^{(1)}_{v}=15$, $g^{s}_{v}=18$, $g^{a}_{v}=18$. Fig. 11(a) shows the resulting triple point for $B_{\perp}=0$. After double-clicking a point in gate space, Fig. 11(b) displays the corresponding current and configuration of the chemical potentials in the right and left QD. The left horizontal lines in (b) depict the chemical potential of the four possible states of the single electron in the left QD. The right horizontal lines depict the chemical potentials of transitions from $N=1\rightarrow 2$, into the orbitally symmetric states |$\ket{S^{v}T^{s}_{0\pm}}$ (blue), |$\ket{T^{v}_{0\pm}S^{s}}$ (green), and the orbitally antisymmetric states, |$\ket{T^{v}_{0\pm}T^{s}_{0\pm}},\ket{S^{v}S^{s}}$ (red), assuming the single electron in the right QD to be in the GS before the transition. Fig. 11(c) and (d) show the corresponting probability vector $\mathrm{P}_{\chi}$ and the rate matrix $W_{\chi\leftarrow\chi^{\prime}}$. Having clicked at the position of the blue cross, we can indeed verify the interpretation of Fig. 3, that the blockade is lifted as soon as the detuning compensates the orbital splitting and the antisymmetric states become accessible. In the probability vector we can see that the DQD is in the $(1,1)$ configuration most of the time, i.e. we are limited by the interdot tunnel rate, just as in our experimental data.

### v2

@@ Abstract
To electrostatically define a DQD, we follow previous works[22, 23, 24, 25, 50, 51, 26] and fabricate heterostructures consisting of a BLG crystal encapsulated within two layers of hexagonal boron nitride (with a thickness of $\approx 30$ nm each) [52, 53]. The heterostructure is placed on a graphite flake which acts as a back gate (BG) [54]. Split gates (SGs) consisting of $\approx 25$ nm Cr/Au are deposited on top, forming a $\approx 100$ nm wide channel. An additional $\approx 75$ nm thick layer of metallic Cr/Au finger gates (FGs) is then fabricated across the channel. The top gate layers are separated by a 15 nm thick layer of atomic layer deposited Al2O3. A schematic top view of a finished sample is shown in Fig. 1(a). All measurements are perfomed in a dilution refrigerator with a base temperature below 30 mK. Voltages of opposite sign applied to the SGs and BG create an electric displacement field perpendicular to the plane of the BLG, opening a band gap and allowing the Fermi energy to be tuned into the gap [54, 35]. This leaves a narrow conductive channel, defined by the SGs, connecting the source and drain reservoirs. Voltages applied to two adjacent FGs ($V_{\mathrm{L}},V_{\mathrm{R}}$) are used to locally invert the polarity of the channel, creating a DQD and giving rise to tunnel barriers where the Fermi energy lies within the band gap, as illustrated in Fig. 1(b). At a finite bias voltage ($V_{\mathrm{SD}}$), we measure the conduction through the channel as a function of $V_{\mathrm{L}}$ and $V_{\mathrm{R}}$, as shown in Fig. 1(c). For $V_{\mathrm{L}}>4.2$ V and $V_{\mathrm{R}}>4.05$ V, we observe the formation of an electron-electron DQD, which is incrementally filled for increasingly positive FG voltages. White numbers $(N_{\mathrm{L}},N_{\mathrm{R}})$, indicate the charge occupation in the left and right QD, respectively. Note that for smaller $V_{\mathrm{L}}$ and $V_{\mathrm{R}}$, a hole QD is formed between the two electron QDs, which appears as curved Coulomb resonances [55, 25]. For larger $V_{\mathrm{L}}$, $V_{\mathrm{R}}$, both the interdot tunnel coupling and the tunnel coupling to the leads electron-electron DQD increase significantly [28]. In particular, the tunnel barrier to the right lead is weakened already at $V_{\mathrm{L}}>4.4$ V, $V_{\mathrm{R}}>4$ V, leading to a strongly pronounced co-tunneling line.
@@
To understand the observed resonant features and the blockade effects, we discuss possible transitions between single- and two-particle states in a DQD when including both symmetric and antisymmetric orbital states. For the investigated energy range we focus on 16 relevant two-particle states, which we denote as spin ($s$) and valley ($v$) singlet ($S$) and triplet ($T_{0},T_{+},T_{-}$) states. For example, |$\ket{T^{s}_{-}S^{v}}$ represents a fully polarized spin triplet and a valley singlet state. The notation $0\pm$ encompasses all triplet states, namely the fully polarized (+,+,-+ , -) and unpolarized (00) triplet states. The relevant states group into 6 orbitally symmetric and 10 orbitally antisymmetric states, separated by the orbital splitting $\Delta_{\mathrm{Orb}}$ [57, 58]. The orbitally symmetric states $\phi_{s}$ are further split by short-range electron-electron interactions into a valley singlet – spin triplet (|$\ket{T^{s}_{0\pm}S^{v}}$) ground state (GS) and valley triplet – spin singlet (|$\ket{S^{s}T^{v}_{0\pm}}$) excited states (ESs) with a separation of $\delta_{2}$ [57, 59, 58, 60, 61], as shown in Fig. 3(a). The next higher orbital state is antisymmetric $\phi_{a}$, featuring spin triplet - valley triplet states (|$\ket{T^{s}_{0\pm}T^{v}_{0\pm}}$) and a spin singlet - valley singlet state (|$\ket{S^{s}S^{v}}$), which are separated by the Kane-Mele spin-orbit coupling $\Delta_{\mathrm{SO}}\approx 65~{}\mu$eV [62, 26, 25]. For the $(1,1)$ states, we assume a negligible mixing between the two QDs due to a small interdot tunnel coupling and treat them as independent single particle states in each QD.
First, we focus at the $(1,1)\rightarrow(2,0)$ charge transition at $V_{\mathrm{SD}}<0$ V. For simplicity, we initially only consider transitions from the $(1,1)$ GS to the 16 different two-particle states. Fig. 3(a) displays possible transitions from the $(1,1)$ GS, which is the |$\ket{K^{-}\uparrow,\,K^{-}\uparrow}$ state for $B_{\perp}\geq 0$ [62, 26, 25], into spin triplet - valley singlet, |$\ket{T^{s}_{0\pm}S^{v}}$ (green), spin singlet - valley triplet, |$\ket{S^{s}T^{v}_{0\pm}}$ (blue) and spin triplet - valley triplet states, |$\ket{T^{s}_{0\pm}T^{v}_{0\pm}}$ (red) of the symmetric and antisymmetric two-particle states in the $(2,0)$ configuration at $B_{\perp}=0$. Increasing the perpendicular magnetic field shifts all states according to their spin and valley magnetic moments [57], which is illustrated in Fig. 3(b). The condition for a charge transition to resonantly occur is that the detuning energy of the DQD needs to compensate the energy difference between the electrochemical potentials of the $(1,1)$ and $(2,0)$ states, given by
@@
with the Bohr magneton, $\mu_{B}$, and the valley g-factor of the single particle orbital, $g^{(1)}_{v}$. For $B_{\perp}>B_{\mathrm{TS}}$, where the two-particle GS changes to |$\ket{S^{s}T^{v}_{-}}$, the GS to GS transition is |$\ket{K^{-}\uparrow,\,K^{-}\uparrow}\rightarrow\ket{S^{s}T^{v}_{-}}$, which shifts as
@@
Importantly, Fig. 3(c) illustrates that we expect to observe a magnetic field tunable valley and spin blockade in the $(1,1)\rightarrow(2,0)$ bias triangle, as indicated by the green and blue colored regions, respectively. For $B_{\perp}<B_{\mathrm{TS}}$, the $(1,1)$ GS to $(2,0)$ GS transition |$\ket{K^{-}\uparrow,\,K^{-}\uparrow}\rightarrow\ket{T^{s}_{-}S^{v}}$ (green region) requires a valley flip. For increasing detuning energy $\varepsilon$, transitions to the excited states |$\ket{S^{s}T^{v}_{0\pm}}$ become energetically accessible (blue arrows in Fig. 3(a)). These transitions are spin blocked. However, the valley blocked transition to the ground state remains energetically available. Since the valley relaxation rate is expected to be faster than the spin relaxation rate  [63, 64], valley relaxation is the process limiting the blockade. Valley blockade (VB) is therefore only lifted if the antisymmetric orbital states are accessible, limiting the detuning energy range of the valley blocked region to
@@
The position of the zero detuning energy in the gate space changes slightly during the measurement, firstly due to slow electrostatic noise in the sample, and secondly due to the GS to GS transition shifting with the magnetic field. Therefore, using magnetic field dependent measurements of the bias triangles (see Fig. 2), the detuning traces are shifted such that $\varepsilon=0$ is always at the same position, as indicated by the white dashed line in Fig. 4(a). Our simulation is able to reproduce many features in the data using $\Delta_{\mathrm{Orb}}=0.7$ meV, $\delta_{2}=0.34$ meV, $g^{(1)}_{v}=15$, $g^{s}_{v}=18$, $g^{a}_{v}=19$, which is in agreement with other measurements on QDs of similar dimension [37, 25, 26]. Both simulation and measurements show the transition from valley blockade (marked by the star) to spin blockade (marked by the circle), as can be seen by comparing Figs. 4(a) and 4(b). We identify prominent resonances labeled a - d in the data  [^1] The software of the simulation allows to interactively inspect any position in gate space, where it illustrates the alignment of the chemical potentials in the left and right QD., with resonance a corresponding to a transition from the $(1,1)$ GS to the valley unpolarized triplet state in the antisymmetric orbital, |$\ket{K^{-}\uparrow,\,K^{-}\uparrow}\rightarrow\ket{T^{s}_{-}T^{v}_{0}}$, which requires a valley flip. Note that this transition is pronounced distinctly, despite the necessary valley flip. This is likely due to the different distribution of the antisymmetrical orbital state (compared to the symmetric orbital state) in k-space, leading to a non-zero overlap between the $(1,1)$ |$\ket{K^{-}K^{-}}$ state and the anti-symmetric |$\ket{T^{v}_{0}}$, which leads to an increased tunnel rate compared to transitions targeting the symmetric |$\ket{T^{v}_{0}}$. Resonance b corresponds to the transition that first lifts the valley and spin blockade, |$\ket{K^{-}\uparrow,\,K^{-}\uparrow}\rightarrow\ket{T^{s}_{-}T^{v}_{-}}$, and is strongly pronounced in both data and simulation, as it is the transition from the (1,1) GS to the antisymmetric orbital state that requires the lowest detuning energy where both spin and valley blockade are lifted. Resonance c involves a $(1,1)$ ES and requires a valley flip, |$\ket{K^{-}\uparrow,\,K^{-}\downarrow}\rightarrow\ket{S^{s}T^{v}_{0}}$. Finally, different valley blocked transitions contribute to resonance d, |$\ket{K^{-}\uparrow,\,K^{+}\downarrow}\rightarrow\ket{T^{s}_{0}T^{v}_{-}}$, |$\ket{K^{-}\downarrow,\,K^{+}\downarrow}\rightarrow\ket{T^{s}_{+}T^{v}_{-}}$, |$\ket{K^{-}\downarrow,\,K^{+}\uparrow}\rightarrow\ket{T^{s}_{0}T^{v}_{-}}$, which is therefore rather broadened. Some of the resonant features do not shift linearly with $B_{\perp}$, which has been observed before in similar measurements and can be caused by the magnetic field slightly changing the effective confinement potential and thereby changing the wavefunction dependent valley g-factors. Additionally, we observe a dip in conductance at $B_{\perp}\approx\pm 0.05$ and 0.10.10.1~{}0.1T. Similar effects can be observed in Fig. 10 and in equivalent measurements of BLG DQDs in the $(1,2)\leftrightarrow(0,3)$ charge configuration reported in Ref. [34], but have not yet been understood comprehensively.
@@
Following the same procedure as presented in the context of Fig. 3(c), we are able to plot the $B_{\perp}$ dependency of the expected resonant features as a function of detuning and perpendicular magnetic field, shown in Fig. 5(a). Here, the change in the two particle GS does not affect the detuning dependency of the transitions, because it also affects the onset of the bias triangle (defining $\varepsilon=0$), see also equation (The role of antisymmetric orbitals and electron-electron interactions on the two-particle spin and valley blockade in graphene double quantum dots). Therefore, Fig. 5(a) basically shows the magnetic field dispersion of the $(1,1)$ states with the energetically lowest state set to zero. For $B_{\mathrm{TS}}>B_{\perp}>0$, we expect an increasingly large valley blocked region (green region in Fig. 5(a)) since the two particle GS is a valley singlet, |$\ket{T^{s}_{-}S^{v}}$, while the $(1,1)$ GS is valley polarized, |$\ket{K^{-}\uparrow,K^{-}\uparrow}$, already at small $B_{\perp}\neq 0~{}$T. The blockade is lifted as soon as the valley unpolarized $(1,1)$ states are available, which occurs at
@@
However, transport is still limited by valley blockade since the measured average tunnel current includes many charge cycles through the DQD. Thus, after a few cycles, |$\ket{K^{-}\uparrow,K^{-}\uparrow}$ will be occupied again by chance, reestablishing the valley blockade. For $B_{\perp}>B_{\mathrm{TS}}$, the two-particle GS becomes a spin singlet - valley triplet, |$\ket{S^{s}T^{v}_{-}}$, immediately lifting the valley blockade. Instead, a small spin blockade (blue region in Fig. 5(a)) appears close to zero detuning since the $(1,1)$ GS is spin polarized. Black arrows in Fig. 5(a) indicate the spin orientation of the three lowest $(1,1)$ states. The spin blockade is lifted as soon as the first spin unpolarized state is available
@@
We experimentally probe the $(2,0)\rightarrow(1,1)$ charge transition by measuring magnetic field dependent detuning cuts through the bias triangle, similar to Fig. 4(a), but for positive bias. These are presented in Fig. 5(b). Using the same parameters as for negative bias direction (see Fig. 4(b)), our simulation is able to reproduce the main features of the experimental data. Interestingly, both data and simulation show that the valley blocked region (star) extends beyond $B_{\mathrm{TS}}\approx 0.4$ T, which is in contrast to Fig. 5(a). Furthermore, the simulation predicts that the spin blocked region (circle) appears along the entire magnetic field range, not only for $B_{\perp}>B_{\mathrm{TS}}$. The spin blocked region predicted by the simulation is barely visible in the data, it can only be surmised close to zero detuning for $B_{\perp}\approx 0.05-0.15$ T. However, this can be caused by difficulties in estimating the onset of the bias triangle due to significant co-tunneling, see white dashed lines in Fig. 2(b).
@@ I.2 Rate equation based transport simulation
The software further allows to interactively inspect the charge stability diagrams, as illustrated by Fig. 11, where the DQD parameters in the simulation were set to $\Delta_{\mathrm{Orb}}=0.575$ meV, $\delta_{2}=0.2$ meV, $\delta_{1}=0$, $g^{(1)}_{v}=15$, $g^{s}_{v}=18$, $g^{a}_{v}=18$. Fig. 11(a) shows the resulting triple point for $B_{\perp}=0$. After double-clicking a point in gate space, Fig. 11(b) displays the corresponding current and configuration of the chemical potentials in the right and left QD. The left horizontal lines in (b) depict the chemical potential of the four possible states of the single electron in the left QD. The right horizontal lines depict the chemical potentials of transitions from $N=1\rightarrow 2$, into the orbitally symmetric states |$\ket{S^{v}T^{s}_{0\pm}}$ (blue), |$\ket{T^{v}_{0\pm}S^{s}}$ (green), and the orbitally antisymmetric states, |$\ket{T^{v}_{0\pm}T^{s}_{0\pm}},\ket{S^{v}S^{s}}$ (red), assuming the single electron in the right QD to be in the GS before the transition. Fig. 11(c) and (d) show the corresponding probability vector $\mathrm{P}_{\chi}$ and the rate matrix $W_{\chi\leftarrow\chi^{\prime}}$. Having clicked at the position of the blue cross, we can indeed verify the interpretation of Fig. 3, that the blockade is lifted as soon as the detuning compensates the orbital splitting and the antisymmetric states become accessible. In the probability vector we can see that the DQD is in the $(1,1)$ configuration most of the time, i.e. we are limited by the interdot tunnel rate, just as in our experimental data.